OpenStudy (anonymous):

If an edge of a square is increased by 10%,then by how much will its surface area be increased?

OpenStudy (anonymous):

increasing a certain thing 'to' or 'by' do u understand the differnce in both of them ?????

OpenStudy (anonymous):

tell me asap

OpenStudy (anonymous):

by

OpenStudy (anonymous):

'by' means certain fraction of total is increased or decreased here in this case it is increased

OpenStudy (anonymous):

it means here u hav to increase or add the that fraction (10/100) to the initial edge of square

OpenStudy (anonymous):

that is 10/100 = 0.1 (fraction of total )

OpenStudy (anonymous):

do u knw the formula of surface area of square?

OpenStudy (anonymous):

6(s*s)

OpenStudy (anonymous):

only s*s

OpenStudy (anonymous):

k gud u knw that

OpenStudy (anonymous):

now see let us suppose we hav the edge of sq. initially be 1 ..then by the defination of 'by' that i hav told u above u can add fraction (0.1) to 1 that is nw the final edge length is 1.1

OpenStudy (anonymous):

follow till here ?

OpenStudy (anonymous):

k

OpenStudy (anonymous):

now tell me what is the surface area of sq. with new edge length 1.1????????

OpenStudy (anonymous):

m singh plss reply??????????/

OpenStudy (anonymous):

wait i m understanding vat u did

OpenStudy (zehanz):

Area of square with edge a is a². Edge of new square is a+10% of a = a + 0.1a = 1.1a. Area of new square is (1.1a)²=1.21a² = a² + 0.21a² = a² + 21% of a². So it is 21 % more.

OpenStudy (anonymous):

@ZeHanz PLS. DON'T GIVE DIRECT ANSWER

OpenStudy (zehanz):

@manishsatywali : Just thought you had done enough to explain. Sometimes it is better to write the whole thing down, to show the steps. If I had done this right away, I would agree with you ;)

OpenStudy (anonymous):

(1.1a)^2

OpenStudy (anonymous):

bt i also encourage the way u deliver that in steps...........gd to see that

OpenStudy (anonymous):

i think u both have well explained @manishsatywali and @ZeHanz

OpenStudy (anonymous):

now have you got the idea...........i mean u reach the final answer any difficulty with the method??????????

OpenStudy (anonymous):

no of course thank u @manishsatywali and @ZeHanz

OpenStudy (anonymous):

@ZeHanz .........our combo helped @msingh......it's the best we can make out of openstudy

OpenStudy (zehanz):

@manishsatywali: thank you!